OpenStudy (anonymous):

A particle starts from rest at the point two units from the origin and moves along a flat track with a constant positive acceleration for time t > or = 0. Which of the following could be the graph of the position s(t) of the particle relative to the origin as a function of time t? and why????

OpenStudy (anonymous):



OpenStudy (anonymous):

Use the formulas eg v= u + at, s=ut + at^2/2 to estimate the graph

OpenStudy (anonymous):

I dont even know what those formulas are...

OpenStudy (anonymous):

OK. forget the math, just think about it. Assume distance is represented along the horizontal axis and time vertical, where is the particle at time t = 0?

OpenStudy (anonymous):

It is  ?

OpenStudy (anonymous):

I mean is it?? *

OpenStudy (anonymous):

am i right or wrong?